# MambaIO: Global-Coordinate Inertial Odometry for Pedestrians via Multi-Scale Frequency-Decoupled Modeling

### v1

@@ III-C MambaIO
Motivated by the analyses in the preceding sections, which indicate that the global coordinate frame provides superior feature representations for pedestrian IO, we further investigate strategies to enhance localization accuracy within this frame. We observe that end-to-end learning directly on raw IMU measurements is susceptible to interference from mixed signal components, including high-frequency noise and redundant local motions, which impedes the model’s ability to focus on salient motion cues. To address this, we propose MambaIO: a novel architecture that first decomposes IMU signals in the global frame into multi-scale frequency components using a Laplacian pyramid, separating the low-frequency component, which primarily encodes global motion trends, from the high-frequency component, which captures local dynamic details. Subsequently, a multi-path convolutional module, specialized for local modeling, processes the high-frequency component, while a Mamba module, adept at long-range sequential modeling, handles the low-frequency component. The outputs from both pathways are then fused to enable precise motion state inference. The overall architecture of MambaIO draws inspiration from the design principles of ConvNeXt, and its schematic diagram is presented in Fig. 6.
@@ III-C1 Laplacian Pyramid Decomposition Module
Existing studies suggest that convolutional operations are inherently effective for capturing local spatiotemporal features, corresponding to high-frequency signal components, whereas attention mechanisms or state space models (e.g., Mamba) excel at modeling long-range dependencies associated with low-frequency trends. Therefore, explicitly decomposing IMU signals by frequency and designing dedicated modeling pathways for different bands can substantially improve the specificity and efficiency of feature learning.
The classical Laplacian pyramid generates multi-resolution low-frequency subbands through successive low-pass filtering and downsampling, followed by upsampling and residual computation to extract high-frequency details at each level. However, directly integrating external frequency decomposition tools, such as FFT-based filters, into deep learning frameworks often results in non-differentiable operations, training instability, or inefficient inference. To circumvent these issues while preserving the essence of multi-scale decomposition, we design a lightweight, differentiable approximation using native PyTorch operations.
Specifically, given the input feature $X\in\mathbb{R}^{N\times L}$ in the global coordinate frame, where $N=6$ denotes the three-axis acceleration and three-axis angular velocity, and $L=200$ is the number of samples in the temporal window, we first employ depthwise convolution to approximate low-pass filtering. The convolution kernel weights are fixed to a normalized averaging kernel $\mathbf{K}_{\text{avg}}\in\mathbb{R}^{k}$, satisfying $\sum_{i}\mathbf{K}_{\text{avg}}(i)=1$, and a stride $s=2$ is applied to compress the sequence length, thereby jointly approximating low-pass filtering and downsampling. In our experiments, we set the kernel size $k=5$, yielding the downsampled low-frequency feature:
$X_{ld}=\text{DWConv}_{\text{avg}}(X)\in\mathbb{R}^{N\times\lfloor L/s\rfloor}$ (4)
@@
$X_{\text{high}}=X-X_{\text{low}}\in\mathbb{R}^{N\times L}$ (6)
@@ III-C2 Multi-Path Convolution Module
$X_{0}$ $=\text{DWConv}_{k=1}(X_{\text{high}})\in\mathbb{R}^{N\times L}$ (7)
$X_{1}$ $=\text{DWConv}_{k=3}(X_{\text{high}})\in\mathbb{R}^{N\times L}$
$X_{2}$ $=\text{DWConv}_{k=7}(X_{\text{high}})\in\mathbb{R}^{N\times L}$
$X_{\text{concat}}$ $=\text{Concat}(X_{0},X_{1},X_{2})\in\mathbb{R}^{3N\times L}$
@@
Following the design principles of RepViT, we enhance cross-channel interaction by inserting a lightweight Squeeze-and-Excitation (SE) block after concatenation, followed by a pointwise convolution for channel fusion, yielding the final output of the high-frequency branch:
$X_{\text{MPC}}=\text{Conv}\bigl(\text{SEBlock}(X_{\text{concat}})\bigr)\in\mathbb{R}^{N\times L}$ (8)
@@ III-C3 Mamba Module
State Space Models (SSM), inspired by continuous dynamical systems, have been widely adopted to model long-range dependencies in sequential data, achieving strong performance in computer vision and sequential modeling tasks. Mamba extends the classical SSM by incorporating an input-dependent selective mechanism, enabling linear-complexity modeling of global context while effectively focusing on salient information.
@@
$\dot{h}(t)$ $=Ah(t)+Bx(t)$ (9)
$y(t)$ $=Ch(t)$ (10)
where $x(t)\in\mathbb{R}^{S\times D}$ denotes the input sequence (with $S$ as the sequence length and $D$ as the number of channels), $h(t)\in\mathbb{R}^{H}$ is the hidden state of size $H$, $y(t)\in\mathbb{R}^{S\times D}$ is the output, and $A\in\mathbb{R}^{H\times H}$, $B\in\mathbb{R}^{H\times D}$, $C\in\mathbb{R}^{D\times H}$ are learnable parameters.
Discretization: For discrete sequence processing, the system is typically discretized using Zero-Order Hold. Given a sampling interval $\Delta$, the discretized parameters are:
$\tilde{A}$ $=\exp(\Delta A)$ (11)
$\tilde{B}$ $=(\Delta A)^{-1}(\exp(\Delta A)-I)\cdot(\Delta B)$ (12)
@@
Mamba enhances this framework with an input-dependent selection mechanism that improves its ability to attend to relevant contextual cues. As the internal mechanics of Mamba are not central to this work, we omit further details. We implement the Mamba module following the MambaVision construction, as illustrated in Fig. 7. Based on the State Space Discretization (SSD) procedure, the computation in the Mamba module can be abstracted as:
$X_{1}$ $=\text{SSD}\bigl(\sigma(\text{Conv}(\text{Linear}(X_{\text{low}})))\bigr)$ (15)
$X_{2}$ $=\sigma(\text{Conv}(\text{Linear}(X_{\text{low}})))$ (16)
$X_{\text{Mamba}}$ $=\text{Linear}\bigl(\text{Concat}(X_{1},X_{2})\bigr)$ (17)
@@
This module captures global motion dependencies from the low-frequency component $X_{\text{low}}$ with $\mathcal{O}(L)$ linear complexity, complementing the high-frequency MPC module. Prior work indicates that adding attention mechanisms in deeper layers can further improve modeling accuracy; therefore, a standard attention module is appended after the final Mamba block to enhance the representational weight of critical frames.
@@
$X_{\text{out}}=\text{Conv}(\text{Concat}(X_{\text{MPC}},X_{\text{Mamba}}))\in\mathbb{R}^{N\times L}$ (18)
@@ IV-A1 Datasets
We conduct experiments on six publicly available pedestrian inertial odometry datasets: RIDI, RoNIN, RNIN, OxIOD, TLIO, and IMUNet. These datasets cover diverse indoor and outdoor environments and include a wide range of motion patterns, such as walking, running, and stair climbing, providing a comprehensive basis for evaluating model generalization and robustness.
@@ IV-A2 Baseline Methods
Considering that learning-based approaches substantially outperform traditional physics-driven methods in localization accuracy, we select representative deep learning models as baselines. These include CNN-based methods: RoNIN-ResNet, IMUNet, DeepLite, DeepILS, and the network proposed in TLIO; attention- or LSTM-based methods: RoNIN-LSTM, CTIN, iMOT, and SPBIT; and hybrid architectures, such as the network proposed in RNIN. In addition, following the practices in IMUNet and DeepLite, we include general-purpose networks that have demonstrated strong performance in other domains, including RepViT, MobileNetv3, and MobileNetV4, to enable fair cross-architecture comparison.
@@ IV-A3 Evaluation Metrics
We adopt two widely used metrics in inertial odometry: Absolute Trajectory Error (ATE) and Relative Trajectory Error (RTE). ATE computes the root-mean-square error between the estimated trajectory and the ground-truth trajectory after global alignment, reflecting overall localization accuracy. RTE measures trajectory error within a fixed-length sliding window, evaluating short-term consistency and drift control.
@@ IV-A4 Training Details
The model employs four hierarchical feature extraction stages with channel dimensions [64, 128, 256, 512]. We optimize the network using Adam with an initial learning rate of $10^{-4}$ and terminate training early if the learning rate decays to $10^{-6}$. The maximum number of training epochs is set to 40, based on validation set observations indicating that most models begin to overfit beyond this point. All experiments are conducted on five NVIDIA RTX 3090 GPUs using CUDA 12.2 and PyTorch 2.5.1.
@@ IV-B Coordinate Frame Representation Comparison
Figure 8 presents the ATE and RTE results of all baseline methods when trained and evaluated under the body-fixed and global coordinate frames separately. It is evident that, in the body-fixed frame, most models struggle to learn effective motion representations, resulting in substantially lower localization accuracy compared to their performance in the global frame. This finding is consistent with the trajectory visualization analysis in  III-B, which shows that the global coordinate frame provides a more stable and discriminative motion representation, thereby markedly enhancing model performance. Detailed comparative results across additional datasets are provided in the Appendix.
@@ IV-C Algorithm Comparison
Table 1 summarizes the quantitative results of all methods on the six public pedestrian IO datasets. The experiments demonstrate that the proposed MambaIO consistently outperforms existing state-of-the-art approaches across most scenarios. For example, on the RoNIN dataset, MambaIO reduces ATE by approximately x% and RTE by about x% compared to RoNIN-ResNet. On the largest dataset, TLIO, MambaIO further improves over the latest Transformer-based inertial localization method, reducing ATE by x%–x% and RTE by x%–x%. These results quantitatively confirm the significant gains in localization accuracy achieved by our method.
To provide a visual assessment of localization performance, we present representative trajectories from each dataset (Fig. 9). For simple straight-line or gently curved trajectories (Fig. 9(a)–(c)), both RoNIN-ResNet and MambaIO closely follow the ground-truth paths. In contrast, for trajectories with multiple turns or complex dynamics (Fig. 9(d)–(f)), RoNIN-ResNet exhibits noticeable deviations, whereas MambaIO maintains high accuracy. This improvement is attributed to the Laplacian pyramid’s multi-scale decomposition of IMU signals, which allows the model to jointly capture local details and global trends, resulting in a more robust and precise motion representation.
@@ IV-D Ablation Study
To evaluate the effectiveness of the proposed components, we construct two ablation variants: (1) Conv-only, which processes raw IMU signals using only the multi-path convolution module; and (2) Mamba-only, which processes raw IMU signals using only the Mamba module. Table 2 presents the performance of these variants, the full MambaIO model, and RoNIN-ResNet across the six datasets.
The results indicate that while Conv-only achieves basic localization capability, its accuracy is substantially lower than that of the full MambaIO, suggesting that local modeling alone is insufficient to capture the complex multi-scale dynamics in IMU signals. Similarly, Mamba-only underperforms, highlighting the limited ability of purely global modeling to represent high-frequency local details such as gait transients or turning jitter. Notably, both ablation variants still outperform RoNIN-ResNet, confirming the intrinsic effectiveness of the designed modules. The full MambaIO achieves the best performance across all datasets, significantly surpassing current state-of-the-art methods, which strongly validates the complementary benefits of the high/low-frequency decomposition strategy and the dual-path collaborative architecture.
In summary, the ablation study not only confirms the necessity of each component but also underscores the critical role of multi-scale feature disentanglement and joint local–global modeling in enhancing pedestrian inertial localization accuracy.

### v2

@@ III-C MambaIO
The preceding analysis and experiments demonstrate that the global frame yields motion representations with higher observability. However, an effective method for accurately modeling the implicit pose transformations and motion information embedded in IMU measurements remains lacking. We argue that IMU signals contain multi-scale motion characteristics mixed with noise and other interfering components, which collectively hinder a model’s ability to focus on the most relevant information. To address this challenge, we propose MambaIO, which first decomposes IMU measurements in the frequency domain and then extracts informative features tailored to the distinct characteristics of each frequency band. The overall architecture is illustrated in Fig. 6.
@@ III-C1 Laplacian Pyramid Decomposition Module
Existing studies suggest that convolutional operations are inherently effective at capturing local spatiotemporal features, which correspond to high-frequency signal components, whereas attention mechanisms or state space models (e.g., Mamba) excel at modeling long-range dependencies associated with low-frequency trends. Therefore, explicitly decomposing IMU signals by frequency and designing dedicated modeling pathways for different bands can substantially improve the specificity and efficiency of feature learning.
Figure 5: Trajectory predictions by RoNIN ResNet on the RoNIN dataset under Body (blue) and Global (orange) coordinate representations. The Global-frame trajectory aligns more closely with the ground truth (green), indicating improved localization accuracy.
Classical frequency decomposition methods, such as the Laplacian pyramid, generate multi-resolution low-frequency subbands through successive low-pass filtering and downsampling, followed by upsampling and residual computation to extract high-frequency details at each level. The visualization of this process is shown in Fig. 6(b). However, directly integrating external frequency decomposition tools—such as FFT-based filters—into deep learning frameworks often introduces non-differentiable operations, training instability, or inefficient inference. To circumvent these issues while preserving the essence of multi-scale decomposition, we design a lightweight, fully differentiable approximation using native PyTorch operations.
Specifically, given the input feature $X\in\mathbb{R}^{N\times L}$ in the global frame, where $N=6$ denotes the three-axis acceleration and three-axis angular velocity, and $L=200$ is the number of samples in the temporal window, we first employ depthwise convolution to approximate low-pass filtering. The convolution kernel weights are fixed to a normalized averaging kernel $\mathbf{K}_{\text{avg}}\in\mathbb{R}^{k}$, satisfying $\sum_{i}\mathbf{K}_{\text{avg}}(i)=1$, and a stride $s=2$ is applied to compress the sequence length, thereby jointly approximating low-pass filtering and downsampling. In our experiments, we set the kernel size $k=5$, yielding the downsampled low-frequency feature:
$X_{ld}=\text{DWConv}_{\text{avg}}(X)\in\mathbb{R}^{N\times\lfloor L/s\rfloor}.$ (4)
@@
$X_{\text{high}}=X-X_{\text{low}}\in\mathbb{R}^{N\times L}.$ (6)
@@ III-C2 Multi-Path Convolution Module
$X_{0}$ $=\text{DWConv}_{k=1}(X_{\text{high}})\in\mathbb{R}^{N\times L},$ (7)
$X_{1}$ $=\text{DWConv}_{k=3}(X_{\text{high}})\in\mathbb{R}^{N\times L},$
$X_{2}$ $=\text{DWConv}_{k=7}(X_{\text{high}})\in\mathbb{R}^{N\times L},$
$X_{\text{concat}}$ $=\text{Concat}(X_{0},X_{1},X_{2})\in\mathbb{R}^{3N\times L},$
@@
We then enhance cross-channel interaction by inserting a lightweight Squeeze-and-Excitation (SE) block after concatenation, followed by a pointwise convolution for channel fusion, yielding the final output of the high-frequency branch:
$X_{\text{MPC}}=\text{Conv}\bigl(\text{SEBlock}(X_{\text{concat}})\bigr)\in\mathbb{R}^{N\times L},$ (8)
@@ III-C3 Mamba Module
State Space Models (SSMs), inspired by continuous dynamical systems, have been widely adopted to model long-range dependencies in sequential data. Mamba extends the classical SSM by incorporating an input-dependent selective mechanism, enabling linear-complexity modeling of global context while effectively focusing on salient information[Mamba].
@@
$\dot{h}(t)$ $=Ah(t)+Bx(t),$ (9)
$y(t)$ $=Ch(t),$ (10)
where $x(t)\in\mathbb{R}^{S\times D}$ denotes the input sequence (with $S$ as the sequence length and $D$ as the number of channels), $h(t)\in\mathbb{R}^{H}$ is the hidden state of dimension $H$, $y(t)\in\mathbb{R}^{S\times D}$ is the output, and $A\in\mathbb{R}^{H\times H}$, $B\in\mathbb{R}^{H\times D}$, $C\in\mathbb{R}^{D\times H}$ are learnable parameters.
Discretization: For discrete sequences, the system is typically discretized using the Zero-Order Hold method. Given a sampling interval $\Delta$, the discretized parameters are:
$\tilde{A}$ $=\exp(\Delta A),$ (11)
$\tilde{B}$ $=(\Delta A)^{-1}(\exp(\Delta A)-I)(\Delta B),$ (12)
@@
Mamba enhances this framework with an input-dependent selection mechanism that improves its ability to attend to relevant contextual cues. Since the internal mechanics of Mamba are not central to this work, we omit further details. As illustrated in Fig. 6 (a), based on the State Space Discretization (SSD) procedure, the computation in the Mamba module can be abstracted as:
$X_{1}$ $=\text{SSD}\bigl(\sigma(\text{Conv}(\text{Linear}(X_{\text{low}})))\bigr),$ (15)
$X_{2}$ $=\sigma(\text{Conv}(\text{Linear}(X_{\text{low}}))),$ (16)
$X_{\text{Mamba}}$ $=\text{Linear}\bigl(\text{Concat}(X_{1},X_{2})\bigr),$ (17)
@@
This module captures global motion dependencies from the low-frequency component $X_{\text{low}}$ with $\mathcal{O}(L)$ linear complexity, complementing the high-frequency MPC module.
@@
$X_{\text{out}}=\text{Conv}(\text{Concat}(X_{\text{MPC}},X_{\text{Mamba}}))\in\mathbb{R}^{N\times L}.$ (18)
TABLE I: Algorithm performance across datasets. Best and second-best results are marked in orange and yellow, respectively.
Dataset RIDI RoNIN RNIN IMUNet OxIOD Average
Metrics (Performance: meters) ATE RTE ATE RTE ATE RTE ATE RTE ATE RTE ATE RTE
Hybrid- based MambaIO (Ours) 1.750 1.996 4.480 3.726 3.002 2.166 5.303 3.882 1.568 0.980 3.221 2.550
RNIN (ISMAR 2021)[RNIN-VIO] 2.771 2.992 6.657 4.079 4.203 2.285 8.400 5.055 3.662 4.337 5.139 3.150
CNN- based RoNIN ResNet (ICRA 2020)[RoNIN] 2.429 2.599 6.125 4.080 3.924 2.621 8.159 4.706 3.498 1.305 4.827 3.062
TLIO (RA-L 2020)[TLIO] 2.240 2.524 5.522 3.799 4.297 3.353 7.490 5.273 2.020 0.969 4.440 3.184
IMUNet (TIM 2024)[IMUNet] 3.201 3.240 6.071 4.043 3.156 2.262 8.676 5.947 1.638 1.043 4.548 3.307
DeepILS (IOT-J 2025)[DeepILS] 1.893 2.363 5.803 3.903 3.675 2.392 7.861 5.585 3.294 1.233 4.505 3.095
DeepLite (IOT-J 2025)[DeepLite] 3.658 3.814 6.363 4.382 3.878 2.892 7.459 5.746 2.048 1.112 4.681 3.589
RepVIT-M0.6 (CVPR 2024)[Repvit] 2.523 2.880 5.923 3.762 3.916 2.503 7.103 5.141 2.859 1.165 4.465 3.090
MobileNetV4-S (arXiv 2024)[MobileNetV4] 2.677 3.141 5.256 3.732 3.768 2.567 7.787 6.661 4.408 1.617 4.779 3.544
Tartan IMU (CVPR 2025)[Tartan-IMU] 2.254 2.581 5.212 3.748 3.169 1.982 6.291 4.492 1.670 0.940 3.719 2.749
Transformer- based iMOT (AAAI 2025)[iMOT] 2.788 3.485 5.967 4.856 6.359 3.229 8.953 6.616 21.802 5.197 9.174 4.677
CTIN (AAAI 2022)[CTIN] 2.325 2.918 4.878 4.063 4.732 2.813 5.589 4.522 4.846 1.563 4.474 3.176
SBIPTVTL (CSCWD 2024)[SBIPTVT] 2.074 2.317 5.300 3.799 3.781 2.269 6.466 4.325 1.732 0.954 3.871 2.733
Improvement MambaIO vs RoNIN ResNet(%) 28.0 23.2 26.9 8.7 23.5 17.4 35.0 17.5 55.2 24.9 33.3 16.7
Figure 7: Comparison of trajectory error for RoNIN ResNet under global and body coordinate frames across multiple datasets.
@@ IV-A1 Datasets
We conduct experiments on five publicly available pedestrian IO datasets: RIDI[RIDI], RoNIN[RoNIN], RNIN[RNIN-VIO], OxIOD[Oxiod], and IMUNet[IMUNet]. These datasets span diverse indoor and outdoor environments and encompass a wide range of motion patterns—including walking, running, and stair climbing—providing a comprehensive basis for evaluating model generalization and robustness.
@@ IV-A2 Baseline Methods
Given that learning-based approaches substantially outperform traditional physics-driven methods in localization accuracy, we select representative deep learning models as baselines. These include CNN-based methods: RoNIN ResNet[RoNIN], IMUNet[IMUNet], DeepLite[DeepLite], DeepILS[DeepILS], and the network proposed in TLIO[TLIO]; attention-based methods: CTIN[CTIN], iMOT[iMOT], and SBIPTVTL[SBIPTVT]; and hybrid architectures, such as the network proposed in RNIN[RNIN-VIO] and Tartan IMU[Tartan-IMU]. In addition, following the practices in IMUNet and DeepLite, we also include general-purpose networks that have demonstrated strong performance in other domains—namely RepViT[Repvit], and MobileNetV4[MobileNetV4]—to enable fair cross-architecture comparison.
@@ IV-A3 Evaluation Metrics
We adopt two widely used metrics in IO: Absolute Trajectory Error (ATE) and Relative Trajectory Error (RTE)[Ionet]. ATE computes the Root Mean Square Error (RMSE) between the estimated trajectory and the ground-truth trajectory after global alignment, reflecting overall localization accuracy. RTE computes the RMSE within a fixed-length sliding window, evaluating short-term consistency and drift control.
@@ IV-A4 Training Details
The model employs four hierarchical feature extraction stages with channel dimensions [64, 128, 256, 512]. We optimize the network using the Adam optimizer with an initial learning rate of $10^{-4}$ and terminate training early if the learning rate decays to $10^{-6}$. The maximum number of training epochs is set to 40. All experiments are conducted on five NVIDIA RTX 3090 GPUs using CUDA 12.2 and PyTorch 2.5.1.
@@ IV-B Coordinate Frame Representation Comparison
Fig. 7 presents the ATE and RTE curves of RoNIN ResNet when evaluated on IMU data represented in the body-fixed versus the global frame. It is evident that, under the body frame, RoNIN ResNet struggle to learn effective motion representations, resulting in substantially degraded localization accuracy compared to the global frame. This finding aligns with the trajectory visualization analysis in Section III-B: the global frame provides more stable and discriminative motion representations, thereby significantly enhancing model performance. Detailed comparisons across additional models are provided in the Appendix.
TABLE II: The ablation study results. The best result per metric is highlighted with orange background. All metrics are in meters.
Dataset Metric Conv-only Mamba-only MambaIO (full)
RoNIN ATE 6.835 6.063 4.480
RTE 4.160 4.000 3.726
RIDI ATE 2.318 2.114 1.750
RTE 2.656 2.419 1.996
IMUNet ATE 6.132 6.366 5.303
RTE 4.207 4.493 3.882
RNIN ATE 6.154 3.907 3.002
RTE 3.223 2.847 2.166
OxIOD ATE 1.844 2.723 1.568
RTE 1.020 1.167 0.980
@@ IV-C Algorithm Comparison
Table I summarizes the quantitative results of all methods on the five public pedestrian IO datasets. The experiments demonstrate that the proposed MambaIO consistently outperforms existing state-of-the-art approaches across most scenarios. For example, on the RoNIN dataset, MambaIO reduces ATE by approximately $26.9\%$ and RTE by about $8.7\%$ compared to RoNIN ResNet. Furthermore, the average results across all five datasets show that MambaIO achieves the highest localization accuracy, with improvements of $33.3\%$ in ATE and $16.7\%$ in RTE over RoNIN ResNet. These results quantitatively confirm the significant gains in localization accuracy achieved by our method.
Figure 8: Visualization of sample trajectories across five datasets, comparing MambaIO with Tartan IMU.
To provide a visual assessment of localization performance, we present representative trajectories from each dataset (Fig. 4). For simple strATEht-line or gently curved trajectories (Fig. 4 (a), (c), (d) and (c)), both Tartan IMU and MambaIO closely follow the ground-truth paths. In contrast, for trajectories with multiple turns or complex dynamics (Fig. 4 (b) and (f)), Tartan IMU exhibits noticeable deviations, whereas MambaIO maintains high accuracy. This improvement is attributed to the Laplacian pyramid’s multi-scale decomposition of IMU signals, which enables the model to jointly capture local details and global trends, resulting in a more robust and precise motion representation.
@@ IV-D Ablation Study
To evaluate the effectiveness of the proposed components, we construct two ablation variants: (1) Conv-only, which processes raw IMU signals using only the multi-path convolution module; and (2) Mamba-only, which processes raw IMU signals using only the Mamba module. Table II presents the performance of these variants and the full MambaIO model across the five datasets.
The results indicate that while Conv-only achieves basic localization capability, its accuracy is substantially lower than that of the full MambaIO, suggesting that local modeling alone is insufficient to capture the complex multi-scale dynamics in IMU signals. Similarly, Mamba-only underperforms, highlighting the limited ability of purely global modeling to represent high-frequency local details such as gait transients or turning jitter. Notably, the full MambaIO achieves the best performance across all datasets, significantly surpassing current state-of-the-art methods, which strongly validates the complementary benefits of the high/low-frequency decomposition strategy and the dual-path collaborative architecture.